if a+b+c=pie and tan1/4(b+c-a)tan1/4(c+a-b)tan1/4(a+b-c)=1,prove that 1+cos a +cos b +cos c=0

Grade:

1 Answers

Monil Shah
11 years ago

Let A=b+c-a,B=c+a-b,C=a+b-c,then 

A+b+c=pie

Divide by 4 both sides

Multiply by (tan) on both sides and solve the equation.